# Benchmarking Attacks on Learning with Errors

### v1

@@ 5.1 Analysis of Results
For the two attacks implemented in the lattice estimator (uSVP and dual hybrid MiTM), we also provide cost estimates from the estimator (commit f18533a) for each benchmark setting. We only present estimates for the Chen Nguyen cost model, which best approximates the fplll implementation of BKZ SVP and should most closely match our experimental results. Estimates can be found in Table X. The Estimator natively supports sparse ternary ${{\mathfrak{B}}^{-}_{h}}$, but we add a functions in nd.py to estimate attack performance on fixed $h$ binomial secrets, ${{\mathfrak{B}}^{\eta}_{h}}$ and ${\mathcal{N}({\sigma})_{h}}$. See Appendix .1 for details. A script to generate these estimates will be included in our open-source codebase.
@@ 6.1 Q-ary Lattice Reduction
When reducing a lattice basis of the form :
@@
the short vectors are not always balanced, contrary to the common assumption in the literature. Figure 2 illustrates the standard deviations of the coefficients of the reduced vectors for $n=m=256$, $q=3329$, and $\omega=1$. It reveals that the BKZ algorithm reduces vectors from right to left, and depending on the block size $\beta$, it terminates at a certain point, leaving a portion of the vector unreduced. However, when $\beta$ is large enough, BKZ successfully reduces all components of the initial basis, producing balanced vectors. This unbalanced profile, initially demonstrated in the Cool & Cruel attack of [43], admits a powerful attack on sparse secrets.
Figure 2: Entry-wise standard deviation of BKZ2-Reduced vectors of a q-ary lattice generated by the rows of $\mathbf{\Lambda}$ as in 6.1, with parameters $n=m=256,q=3329,\omega=1$ with BKZ2 $\beta$ values. The reduced vectors are not balanced.
If $\bf U=(L,R)$ is the unimodular matrix produced by BKZ applied to $\mathbf{\Lambda}$, then the resulting short vectors whose entries are shown in Figure 2 are the rows of
$\mathbf{U\Lambda}=(\omega\mathbf{R},\mathbf{RA}\mod q)$
The left half of Figure 2 reveals that the rows of $\bf R$ are not balanced either. This side of the graph plots $\omega\mathbf{R}$, and we see that the last columns of $\mathbf{R}$ have only small values. This means the BKZ algorithm underutilizes the later rows of the matrix $\mathbf{A}$. Future work should devote more study to this phenomenon and potential ways to exploit it.
6.2 Discrepancy in BKZ timings
Next, we highlight an interesting experimentally observed discrepancy in the predicted and actual BKZ loop time. We ran BKZ2 with $\beta$ ranging from $40$ to $54$ on qary-embedded (Equation 1) ($m\times n$) LWE matrices with $n=512$, $m=712$, $q=3329$, and $\omega=4$ and measured the time it took for BKZ2 to complete one loop. We then computed the predicted loop cycle/time for BKZ with these setting with two enumeration SVP cost models: CheNgu12 [21] and ABLR21 [6] (see Appendix .2 for details). Concrete BKZ loop times and estimated timings are in Table XIII.
@@ 6.2 Discrepancy in BKZ timings
Our experimental results do not closely match either cost model. We observe a sharp increase in the BKZ2 loop time starting around $\beta=49$. For $\beta\geq 52$, BKZ2 takes several days to run, sometimes failing to terminate within our $72$ hour cluster time limit.
This discrepancy may be due to either some implementation issue or there is some discrepancy in the estimates themselves. More rigorous experimentation with BKZ in higher dimensions and $\beta$ is needed to explore this.
@@ 6.3 ML attacks recover secrets with $\leq 3$ cruel bits
This explains why the ML attack only recovers $3$ cruel bit secrets: models struggle to learn modular arithmetic (as first observed in [37]), and secrets with $>3$ cruel bits require models to understand that certain $b$ elements “wrap around” the modulus. Prior work has observed models’ poor performance on modular arithmetic setups [32, 45, 30, 35], and this result confirms the difficult persists. Future improvements to the ML attack could focus on strategies to help models learn modular arithmetic.
@@ 6.4 Effect of bad PRNGs on attack performance
We highlight this issue because even though PRNGs are well-known to be bad for cryptography, this advice might not always be followed. Furthermore, the fact that LCGs are included as the standard RNG in some libraries increases the likelihood that they may be accidentally used. Using LCG-generated $\bf A$ matrices in LWE attack development would make the attack appear exceptionally good, while LCG use in LWE encryption schemes would create significant vulnerabilities. We observe at least one case of a LCG PRNG used in real-world LWE crypto: the C++ rand function is used in the testing functions of the HEEAN library [18]111github.com/snucrypto/HEAAN/blob/master/HEAAN/src/EvaluatorUtils.cpp. While this use poses no imminent danger, we highlight this as a cautionary tale for implementers of Kyber and HE.
@@ 7 Join our LWE Attack Benchmarking Effort
Contributing. We invite contributors to reproduce our results, improve on these methods, and/or implement new LWE attacks. We actively welcome pull requests with new attacks or code improvements. Please document added code, provide proof of testing, and follow a style similar to the rest of the repository. We will also use Github issues to track public bugs. Please provide a clear description of the bug and instructions on how to reproduce the problem. See the CONTRIBUTING file in our codebase for instructions on how to contribute.
We will also maintain a centralized leaderboard of the best performing attacks on LWE along the axes of time, space, and compute. We welcome pull requests with new or improved attacks that outperform our current benchmark. Please include the code and instructions on how to reproduce the results in the pull request. We will update the leaderboard on our associated challenge website accordingly.
@@ 8 Conclusion and Future Work
Optimizing the attack implementations we provide. Although we do our best to fairly compare the four attacks we evaluate, there are inevitable inefficiencies in our implementations. We expect that additional engineering would make these attacks more efficient. For example, making the enumeration SVP in BKZ faster (via a GPU implementation like [46]) would greatly improve the lattice reduction time for all attacks.
Revisiting theoretical estimates with experimental insights. In several instances, theoretical predictions do not match experimental observations. For example, the Estimator over-estimates the number of short vectors needed but under-estimates time needed for successful Dual Hybrid MiTM attacks. We also observe discrepancies between predicted and actual BKZ2 times, as shown in §6.2 and Appendix .2. These indicate a need for examination of theoretical assumptions to ensure they align with real-world behavior.

### v2

@@ 5.1 Analysis of Results
For the two attacks implemented in the lattice estimator (uSVP and dual hybrid MiTM), we also provide cost estimates from the estimator (commit f18533a) for each benchmark setting. We only present estimates for the Chen Nguyen cost model, which best approximates the fplll implementation of BKZ SVP and should most closely match our experimental results. Estimates can be found in Table X. The Estimator natively supports sparse ternary ${{\mathfrak{B}}^{-}_{h}}$, but we add a function in nd.py to estimate attack performance on fixed $h$ binomial secrets, ${{\mathfrak{B}}^{\eta}_{h}}$ and ${\mathcal{N}({\sigma})_{h}}$. See Appendix .1 for details. A script to generate these estimates will be included in our open-source codebase.
@@ 6.1 Q-ary Lattice Reduction
The Cool & Cruel attack paper [44] showed that when reducing a lattice basis of the form:
@@
the short vectors are not always balanced, contrary to the common assumption in the literature. In particular, letting $\mathbf{U\Lambda}=(\omega\mathbf{R},\mathbf{RA}\mod q)$ be the reduced matrix obtained by applying BKZ2.0 to $\mathbf{\Lambda}$, Figure 2 illustrates the standard deviations of the columns of $U\Lambda$ for $n=m=256$, $q=3329$, and $\omega=1$. It reveals that BKZ2.0 reduces vectors from right to left. Depending on the block size $\beta$, it terminates at a certain point, leaving a portion of the vector unreduced.  [44] used this to mount a powerful attack on sparse secrets.
Figure 2: Larger BKZ2.0 block size produces data with a smaller cliff. As block size increases (color tends toward yellow), BKZ2.0 yields rows of $U\Lambda$ with more entries tending towards the reduced value. Figure shows parameters $n=m=256$, $q=3329$, $\omega=1$, and varying block size $\beta$.
We make two observations extending [44]. First, we observe experimentally that as the block size $\beta$ increases, BKZ2.0 tends towards reducing all components of the initial basis, producing balanced vectors, see Figure 2. Furthermore, we see that BKZ2.0 underutilizes the later rows of the matrix $\mathbf{A}$. The left half of Figure 2 reveals that the rows of $\bf R$ are not balanced either. This side of the graph plots $\omega\mathbf{R}$, and we see that the last columns of $\mathbf{R}$ have only small values, indicating that BKZ2.0 has not fully used the later rows of $\mathbf{A}$. Future work should study this phenomenon and ways to exploit it.
6.2 Discrepancy in BKZ2.0 timings
Next, we highlight an interesting experimentally observed discrepancy in the predicted and actual BKZ2.0 loop time. We ran BKZ2.0 with $\beta$ ranging from $40$ to $54$ on q-ary-embedded (Equation 1) ($m\times n$) LWE matrices with $n=512$, $m=712$, $q=3329$, and $\omega=4$ and measured the time it took for BKZ2.0 to complete one loop. We then computed the predicted loop cycle/time for BKZ2.0 with these settings with two enumeration SVP cost models: CheNgu12 [20] and ABLR21 [6] (see Appendix .2 for details). Concrete BKZ2.0 loop times and estimated timings are in Table XIII.
@@
Our experimental results do not closely match either cost model. We observe a sharp increase in the BKZ2.0 loop time starting around $\beta=49$. For $\beta\geq 52$, BKZ2.0 takes several days to run, sometimes failing to terminate within our $72$ hour cluster time limit.
This discrepancy may be due to some implementation issue or to problems with the estimates themselves. We observe that the cost model derived from [20] is a linear curve fit to experimental results of BKZ2.0 runs on $n\leq 250$. (see CheNgu12 function in reduction.py of the lattice estimator). Other cost models, such as that of [6], build on this curve fit. If the [20] cost model does not generalize well beyond $n=250$, this could explain the error. More rigorous experimentation with BKZ2.0 in higher dimensions and block size $\beta$ is needed to understand this.
@@ 6.3 ML attacks recover secrets with $\leq 3$ cruel bits
This explains why the ML attack only recovers $3$ cruel bit secrets: models struggle to learn modular arithmetic (as first observed in [38]), and secrets with $>3$ cruel bits require models to understand that certain $b$ involve sums which “wrap around” the modulus. Prior work has observed models’ poor performance on modular arithmetic setups [32, 46, 30, 36], and this result confirms that the difficulty persists. Future improvements to the ML attack could focus on strategies to help models learn modular arithmetic.
@@ 6.4 Effect of bad PRNGs on attack performance
We highlight this issue because even though PRNGs are well-known to be bad for cryptography, this advice might not always be followed. Furthermore, the fact that LCGs are included as the standard RNG in some libraries increases the likelihood that they may be accidentally used. Using LCG-generated $\bf A$ matrices in LWE attack development would make the attack appear exceptionally good, while LCG use in LWE encryption schemes would create significant vulnerabilities. We observe at least one case of a LCG PRNG used in real-world LWE crypto: the C++ rand function is used in the testing functions of the HEEAN library [17]github.com/snucrypto/HEAAN/blob/master/HEAAN/src/EvaluatorUtils.cpp. While this use poses no imminent danger, we highlight this as a cautionary tale for implementers of Kyber and HE.
@@ 7 Join our LWE Attack Benchmarking Effort
Contributing. We invite contributors to reproduce our results, improve on these methods, and/or implement new LWE attacks. We actively welcome pull requests with new or improved attacks or code improvements. Please include the code and instructions on how to reproduce the results in the pull request. Please document added code, provide proof of testing, and follow a style similar to the rest of the repository. We will also use Github issues to track public bugs. Please provide a clear description of the bug and instructions on how to reproduce the problem. See the CONTRIBUTING file in our codebase for instructions on how to contribute.
We will also maintain a centralized leaderboard of the best performing attacks on LWE along the axes of time, space, and compute. We will update the leaderboard on our associated challenge website accordingly.
@@ 8 Conclusion and Future Work
Optimizing attack implementations. Although we do our best to fairly compare the four attacks we evaluate, there are inevitable inefficiencies in our implementations. We expect that additional engineering would make these attacks more efficient. For example, speeding up the enumeration SVP in BKZ2.0 (via a GPU implementation like [47]) would greatly improve the lattice reduction time for all attacks. Revisiting theoretical estimates with experimental insights. In several instances, theoretical predictions do not match experimental observations. For example, the Estimator over-estimates the number of short vectors needed but under-estimates time needed for successful Dual Hybrid MiTM attacks. We also observe discrepancies between predicted and actual BKZ2.0 times, as shown in §6.2 and Appendix .2. Rigorous work is needed to better understand, and eventually correct, these discrepancies. In particular, future work should ensure theoretical assumptions align with real-world behavior.